Calculate the price elasticity of demand at price P= $8.50, given demand equation Q= 184 - 4P
Answer #1

Answer :- Demand equation : a = 184 - 48 When P = $8.50 .. Q z 184 -4 18:10). 2184-24 = 150 Price elasticity of demand = SO 1